OpenStudy (anonymous):

-1/3x + y = –1 y = 4+ 1/3x

OpenStudy (anonymous):

if you graph this will it be a solution or not a solution?

OpenStudy (anonymous):

-1/3x + y = –1 solve for y

OpenStudy (anonymous):

can i move the numbers around?

OpenStudy (anonymous):

put it in this form y = mx + b

OpenStudy (anonymous):

like- -1/3x+-1=y

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

didn't you move y to the other side? what should the sign be?

OpenStudy (anonymous):

y=-1+ -1/3x ?

OpenStudy (anonymous):

im lost?

OpenStudy (anonymous):

-1/3x + y = –1 +1/3x +1/3x y = 1/3x - 1

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

y = 1/3x - 1 y = 1/3x + 4 now graph them

OpenStudy (anonymous):

thanks,

OpenStudy (anonymous):

did you graph them?

OpenStudy (anonymous):

i am in a second

OpenStudy (anonymous):

so its a solution?

OpenStudy (anonymous):

do the lines cross somewhere?

OpenStudy (anonymous):

on the a.b.c.d. all of the lines are paralell and they are asking if solution or not.

OpenStudy (anonymous):

lines are parallel so they will never cross, so no solution

OpenStudy (anonymous):

dont forget my $10 :p

OpenStudy (anonymous):

idk its like two paraell lines going up and its says is solution-A then same graph but its say no solution-b

OpenStudy (anonymous):

 no solution

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

thats the right graph

OpenStudy (anonymous):

c and d

OpenStudy (anonymous):

so it b not a solution

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

anytime